Do It Review 26-5 Wayne Company is considering a long-term investment project called ZIP. ZIP will require an investment of $Wayne Company is considering a long-term investment project called ZIP. ZIP will require an investment of $121,720. It will have a useful life of 4 years and no salvage value. Annual revenues would increase by $80,100, and annual expenses (excluding depreciation) would increase by $40,100. Wayne uses the straight-line method to compute depreciation expense. The company’s required rate of return is 13%.

Compute the annual rate of return

Anual rate of return ____%

Determine whether the project is acceptable?

Accept/Reject the project
Answer #1

Annual rate of return = Annual Operating Income/Average Investment

= (80,100-40,100-121720/4)/60,860

= 15.72%

Accept, since Rate of return is higher than the minimum required return

Note: Average Investment = (121720+0)/2 = $60,860